OpenStudy (anonymous):

How can 1/5x-2=1/3x+8 be set up as a system of equations?

OpenStudy (anonymous):

5y - 5x = -10 3y - 3x = 24 5y - 5x = -10 3y - 3x = 24 5y + x = -10 3y + x = 24 5y - x = -10 3y - x = 24